A brick with a length width and height of 10.0cm 2.00cm and 5.00 cm. The pressure exerted by the brick on the floor is different when the brick is placed in different ways. If the minimum pressure exerted by the brick on the floor is 100.0Pa what is the forced exerted on the floor in Newton’s?
1
Expert's answer
2020-05-24T03:09:55-0400

Pressure= "\frac{force }{area}"

pressure will be minimum when area will be maximum

here maximum area = 50cm2 =50"\times"10-4 m2


"100=\frac{force}{50\times10^{-4}}"

Force=0.5 N